OpenStudy (anonymous):

find the equation of a tangenet line when y=8^x at x=1

OpenStudy (phi):

write 8 as \( e^{\ln(8)} \) so that your equation is \[ y = e^{\ln(8)x} \] find the slope dy/dx at x=1

OpenStudy (anonymous):

After finding dy/dx then you would plug that into the line slope equation of (y-yo)=m(x-xo) where m=dy/dx. Right? I am learning too.

OpenStudy (phi):

@Merlined yes, with (x0, y0) = (1,8) Here is a plot